OpenStudy (abbycross167):

How would I put this equation on a graph? Equation~y = -x + 3

OpenStudy (anonymous):

Well first you would have to bring y up to 3 because that's the y intercept

OpenStudy (anonymous):

It's also a negative x so the line goes downward from left to right

OpenStudy (anonymous):

Medals?

OpenStudy (abbycross167):

@livias.random give me a second

OpenStudy (abbycross167):

@atreyu6s3x6 can you help me and see if this graph is right?

OpenStudy (anonymous):

close but not quite the b is the y-intercept so it would go up 3 on the y-axis and the slope is -1/1 give it a shot :) if not ill graph it

OpenStudy (anonymous):

Ah of course sorry didn't see the negative sign

OpenStudy (abbycross167):

ok I think I got it, but would you mind graphing it so I can see if I did it right?

OpenStudy (anonymous):

sure give me a sec :) @abbycross167

OpenStudy (anonymous):

So instead: y = -x + 3 Y intercept: when x = 0, y = -0 + 3 = 3 X intercept: when y = 0, 0 = -x + 3 = 3 So we have the two points (0,3) and (3,0) 

OpenStudy (anonymous):

here you go :D

OpenStudy (abbycross167):

Ok, awesome! Thank y'all soo much both of y'all were very helpful... but I don't know who to give the "best response" thing :/

OpenStudy (anonymous):

nice recovery lol @livias.random

OpenStudy (anonymous):

haha dont matter :) @abbycross167 glad to help :P

OpenStudy (anonymous):

Thanks @atreyu6s3x6 :) And don't worry about it @abbycross167, I'm not that bothered lol

OpenStudy (abbycross167):

@atreyu6s3x6 do you think you could help me with another?

OpenStudy (anonymous):

sure tagg me in @abbycross167

OpenStudy (abbycross167):

ok thank you

OpenStudy (anonymous):

no problem :)